solve each equation.
2/y+ 5/2 = 4 - 2/3y

Answer by tommyt3rd(5050)

2/y+ 5/2 = 4 - 2/(3y)
2/y+ 2/(3y)+ 5/2 - 5/2= 4-5/2 - 2/(3y) + 2/(3y)
8/(3y)=3/2
y= (2/3)(8/3)=16/9